OpenStudy (anonymous):

What is the exact value of the expression the square root of 486. - the square root of 24. + the square root of 6.? Simplify if possible. 2the square root of 2. 2the square root of 6. 6the square root of 2. 8the square root of 6.

OpenStudy (anonymous):

\[\sqrt{486} - \sqrt{24} + \sqrt{6}\] \[= \sqrt{2 \times 3 \times 3 \times 3 \times 3 \times 3} - \sqrt{2 \times 2 \times 2 \times 3} + \sqrt{6}\] \[= 9\sqrt{6} - 2\sqrt{6} + \sqrt{6}\] =\[= 8\sqrt{6}\] Ans.8th the square root of 6 @jboy_52 : your question's answered